OpenStudy (johnkim2330):

HELP PLZ How can one thirdx − 2 = one fourthx + 11 be set up as a system of equations?

OpenStudy (solomonzelman):

write the system in number. For the start.

OpenStudy (johnkim2330):

(1/3)x-2=(1/4)x+11 like this? @SolomonZelman

OpenStudy (solomonzelman):

Okay, so now we can start solving it. \(\Large\color{orangered}{ \frac{1}{3}x-2= \frac{1}{4}x+11 }\) \(\Large\color{orangered}{ \frac{1}{3}x-2-11= \frac{1}{4}x }\) \(\Large\color{orangered}{ -2-11= \frac{1}{4}x -\frac{1}{3}x }\) \(\Large\color{orangered}{ -13= \frac{x}{4} -\frac{x}{3} }\) \(\Large\color{orangered}{ -13= \frac{3x}{12} -\frac{4x}{12} }\) \(\Large\color{orangered}{ -13= -\frac{x}{12} }\) \(\Large\color{orangered}{ 156= x }\)

OpenStudy (johnkim2330):

A.3y − x = −6 4y − x = 44 B.3y + x = −6 4y + x = 44 C.3y − 3x = −6 4y − 4x = 44 D.3y + 3x = −6 4y + 4x = 44 These are the answer choices....i got up to y=(1/3)x-2 y=(1/4)x+11 butnow im not sure what to do

OpenStudy (johnkim2330):

@SolomonZelman

OpenStudy (solomonzelman):

(1/3)x-2=(1/4)x+11 (1/3)x-2=y (1/4)x+11=y and then make all coefficients into whole numbers x-6=3y x+44=4y I think this is what they ask

OpenStudy (johnkim2330):

oh ok...thx a lot

OpenStudy (solomonzelman):

anytime.